Assume nominal income is shown below.
Year. Money income CPI. Real income. P. Power
2005. 430220. 1090.6.
2006. 470000. 1130.6
2007. 500900. 1180.3
2008. 540600. 1240.3
I) determine your real income in 2005
ii)compute purchasing power of dollars for the year 2008
iii)the inflation rate between 2005 and 2006
1
Expert's answer
2020-12-15T11:24:59-0500

i)"real income=\frac{nominal income}{price level}=\frac{430220}{1090.6}=394.48"


ii)the change in the exchange rate is equal to the difference in inflation

1240,3-1180,3=60

"\frac{60}{1180.3}=0.0508" or 5.08%


iii) "inflation rate=\frac{1130.6-1090.6}{1090.6}=0.0367"  or 3.67%